OpenStudy (anonymous):

HELP PLEASE Part A: Use the properties of exponents to explain why 18 to the power of 1/5 is called the fifth root of 18. (5 points) complete sentences Part B: The length of a rectangle is 2 units and its width is (square root symbol) 2 unit. Is the area of the rectangle rational or irrational? Justify your answer. (5 points)

OpenStudy (anonymous):

@SolomonZelman @precal

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@YanaSidlinskiy

OpenStudy (anonymous):

@annas

OpenStudy (anonymous):

@JMark

OpenStudy (anonymous):

@SolomonZelman HELP pleaseee

OpenStudy (yanasidlinskiy):

Part B: it is irrational because the formula for a rectangle is length times width. this would be \[2\sqrt{2}\] and a square root in your answer that cannot be reduced is irrational:)

OpenStudy (anonymous):

Thank you I dont understand part a now :(

OpenStudy (yanasidlinskiy):

\[18(^\frac{ 1 }{ 5 })\] is also called the 5th root of 18 because if you multiply it by itself 5 times, you get 18 \[18(^{\frac{ 1 }{ 5 }}) \times 18(^{\frac{ 1 }{ 5 }}) \times 18(^{\frac{ 1 }{ 5 }}) \times 18(^{\frac{ 1 }{ 5 }}) \times 18(^{\frac{ 1 }{ 5 }}) = 18\] by the rules of exponents \(\huge{\rightarrow}\) \[[18(^\frac{ 1 }{ 5 })^{5}]=18^{1} = 18\]

OpenStudy (anonymous):

Ok thank you!!!

OpenStudy (yanasidlinskiy):

You're Welcome:D Hope that helped:D Sorry it took a little while:D

OpenStudy (anonymous):

Yea haha

OpenStudy (yanasidlinskiy):

Well, my apologies:)